A drug is being delivered into a patient's arm at a rate of 10 mL min-¹. The drug is being delivered from a syringe through a 5 cm long needle with an internal diameter of 1mm. If patient's blood pressure is 110mmHg (+15.0 kPa gauge preasure, and ignoring the variation from systolic to diastolic), what must the pressure in thw syringe be? (ndrug= 8.90x10-4 Pa.s

1
Expert's answer
2022-03-10T18:01:37-0500

"P_g=\frac{\Delta V}{\Delta t}\frac{8L\eta}{\pi r^4}=18.9~kPa."